OpenStudy (anonymous):

In a given training week, a swimmer completes1/4 mile on day one, 2/5 mile on day two, and 3/7 mile on day three. How many total miles does the athlete swim in the given week?

OpenStudy (anonymous):

okay, do you know how to find a common denominator?

OpenStudy (anonymous):

1.07 miles

OpenStudy (anonymous):

Its best to set up your equation like\[\frac{ 1 }{ 4 } + \frac{ 2 }{ 5 } + \frac{ 3 }{ 7 }\]

OpenStudy (anonymous):

then to find your common denominator, find the LCM or least common multiple of 4, 5, and 7...

OpenStudy (anonymous):

Can you find that out on your own? Or do you not know how?